# AV-PedAware: Self-Supervised Audio-Visual Fusion for Dynamic Pedestrian Awareness

### v1

@@ Abstract
In this study, we introduce AV-PedAware, a self-supervised audio-visual fusion system designed to improve dynamic pedestrian awareness for robotics applications. Pedestrian awareness is a critical requirement in many robotics applications. However, traditional approaches that rely on cameras and LIDARs to cover multiple views can be expensive and susceptible to issues such as changes in illumination, occlusion, and weather conditions. Our proposed solution replicates human perception for 3D pedestrian detection using low-cost audio and visual fusion. This study represents the first attempt to employ audio-visual fusion to monitor footstep sounds for the purpose of predicting the movements of pedestrians in the vicinity. The system is trained through self-supervised learning based on LIDAR-generated labels, making it a cost-effective alternative to LIDAR-based pedestrian awareness. AV-PedAware achieves comparable results to LIDAR-based systems at a fraction of the cost. By utilizing an attention mechanism, it can handle dynamic lighting and occlusions, overcoming the limitations of traditional LIDAR and camera-based systems. To evaluate our approach’s effectiveness, we collected a new multimodal pedestrian detection dataset and conducted experiments that demonstrate the system’s ability to provide reliable 3D detection results using only audio and visual data, even in extreme visual conditions. We will make our collected dataset and source code available online for the community to encourage further development in the field of robotics perception systems.
@@ I Introduction
Pedestrian awareness is a crucial factor for many autonomous systems, such as delivery robots [1], COBOTs [2], and manipulators arms [3], in ensuring efficient planning and safety in complex and dynamic environments. The primary approach to achieving pedestrian awareness has been using exteroceptive sensors such as cameras, 3D-LIDARs [4], and UWB [5], which provides high-resolution environmental information. However, these sensors have certain limitations, such as vulnerability to lighting conditions and occlusion, high costs, and the need for additional operator hardware.
@@ II Related work
Previous research in audio has largely centered on natural language processing (NLP) and its use as a means of communication with robots [14, 15]. However, more recently, a growing number of researchers have been exploring the potential of multimodal approaches that integrate audio and visual data to detect and classify abnormal sound sources. For example, one study trained a convolutional neural network (CNN) model for detecting and localizing salient sound sources by combining audio and visual features [9, 10]. Several other studies [16, 17, 18, 19, 11, 12, 13] have also focused on localizing speakers in images based on 3D sound to improve target detection accuracy. Additionally, incorporating visual depth and infrared modalities to train audio networks for detecting moving vehicles has been demonstrated in a recent study [6, 7]. These studies highlight the potential of audio-based approaches for salient sound source detection and target localization.
Previous research has primarily focused on recognizing active speakers or detecting loud sounds, such as vehicles. However, this assumption is not applicable to real-life scenarios where people aren’t always talking or driving cars in office hallways. Instead, humans can correlate visual input with distinct footstep sounds while walking to avoid collisions with other pedestrians, making audio an affordable and perceptible modality. Integrating audio and vision can provide unique advantages such as improved object detection, better occlusion handling, and increased robustness against challenging lighting conditions and noise. However, before incorporating audio into robotic perception systems, challenges such as the limited sensing range of audio sensors, interference from multiple audio sources, and a lack of standardized approaches for multi-sensor calibration and synchronization must be addressed.

### v2

@@ Abstract
In this study, we introduce AV-PedAware, a self-supervised audio-visual fusion system designed to improve dynamic pedestrian awareness for robotics applications. Pedestrian awareness is a critical requirement in many robotics applications. However, traditional approaches that rely on cameras and LIDARs to cover multiple views can be expensive and susceptible to issues such as changes in illumination, occlusion, and weather conditions. Our proposed solution replicates human perception for 3D pedestrian detection using low-cost audio and visual fusion. This study represents the first attempt to employ audio-visual fusion to monitor footstep sounds for the purpose of predicting the movements of pedestrians in the vicinity. The system is trained through self-supervised learning based on LIDAR-generated labels, making it a cost-effective alternative to LIDAR-based pedestrian awareness. AV-PedAware achieves comparable results to LIDAR-based systems at a fraction of the cost. By utilizing an attention mechanism, it can handle dynamic lighting and occlusions, overcoming the limitations of traditional LIDAR and camera-based systems. To evaluate our approach’s effectiveness, we collected a new multimodal pedestrian detection dataset and conducted experiments that demonstrate the system’s ability to provide reliable 3D detection results using only audio and visual data, even in extreme visual conditions. We will make our collected dataset and source code available online for the community to encourage further development in the field of robotics perception systems. Code https://github.com/yizhuoyang/AV-PedAware.
@@
publicationid: pubid: This work has been accepted for publication at the 2023 IEEE/RSJ International Conference on Intelligent Robots and Systems (IROS) © 2023 IEEE. Personal use of this material is permitted. However, permission must be obtained from IEEE for all other uses, including reprinting or redistribution, creating derivative works, or reuse of any copyrighted components of this work in other media.
@@ I Introduction
Pedestrian awareness is a crucial factor for many autonomous systems, such as delivery robots [1], COBOTs [2], and manipulators arms [3], in ensuring efficient planning and safety in complex and dynamic environments [4]. The primary approach to achieving pedestrian awareness has been using exteroceptive sensors such as cameras, 3D-LIDARs [5], and UWB [6], which provides high-resolution environmental information. However, these sensors have certain limitations, such as vulnerability to lighting conditions and occlusion, high costs, and the need for additional operator hardware.
@@ II Related work
Previous research in audio has largely centered on natural language processing (NLP) and its use as a means of communication with robots [15, 16]. However, more recently, a growing number of researchers have been exploring the potential of multimodal approaches that integrate audio and visual data to detect [17] and classify abnormal sound sources. For example, one study trained a convolutional neural network (CNN) model [18] for detecting and localizing [19] salient sound sources by combining audio and visual features [10, 11]. Several other studies [20, 21, 22, 23, 12, 13, 14] have also focused on localizing speakers in images based on 3D sound to improve target detection accuracy. Additionally, incorporating visual depth and infrared modalities to train audio networks for detecting moving vehicles has been demonstrated in a recent study [7, 8]. These studies highlight the potential of audio-based approaches for salient sound source detection and target localization [24, 25, 26, 27, 28, 29].
Previous research has primarily focused on recognizing active speakers or detecting loud sounds, such as vehicles. However, this assumption is not applicable to real-life scenarios where people aren’t always talking or driving cars in office hallways. Instead, humans can correlate visual input with distinct footstep sounds while walking to avoid collisions with other pedestrians, making audio an affordable and perceptible modality [30]. Integrating audio and vision can provide unique advantages such as improved object detection [31], better occlusion handling [32], and increased robustness against challenging lighting conditions and noise. However, before incorporating audio into robotic perception systems, challenges such as the limited sensing range of audio sensors, interference from multiple audio sources, and a lack of standardized approaches for multi-sensor calibration [33] and synchronization must be addressed.